# 1.3in1.3in1in1in

### v1

@@ 3.16. Simple elliptic cubic surfaces
  • the residual to the line $p\in E\simeq F(S)$ is given by $-2p$;
@@
Proposition 3.18.
Let $Y$ be a general cubic fourfold, and let $\ell_{1}$ and $\ell_{2}$ be general lines of type $I$ such that $\psi(\ell_{1})=\psi(\ell_{2})=\ell_{2}$. Then $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$ is a 3-dimensional space which cuts $Y$ in a cubic surface of type either $A_{1}A_{5}$, or ${\widetilde{E}}_{6}$.
Proof.
By Lemma 3.5 a general point of $P$ consists of a pair of lines whose linear span intersects $Y$ in a Cayley cubic. Thus, if $\ell_{1}$ and $\ell_{2}$ are as in the hypothesis, then the 3 dimensional linear space $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$ intersects $Y$ in a cubic surface $S$ which is a degeneration of a Cayley cubic and contains $\ell_{2}$ as a triple line on the surface. By the analysis of lines on cubic surfaces, $S$ is either of type $A_{1}A_{5}$, ${\widetilde{E}}_{6}$ or non-normal. As by assumption the lines $\ell_{1}$ and $\ell_{2}$ are not of type $\operatorname{II},$ we can exclude the non-normal cases. ∎
We explore now some connection of our study of singular cubic surfaces with results of Gounelas and Kouvidakis [GK22]. As a consequence, we deduce the existence of a surface of type $A_{1}A_{5}$, which is a fundamental little ingredient for the main theorem.
Let $Y$ be a smooth cubic fourfold. The natural map $\gamma\colon\mathbb{L}\to Y$ from the universal family of lines to the cubic fourfold is a fibration in (2,3)-complete intersection curves. Given a point $y\in Y$, its fiber $\gamma^{-1}(y)$ consists of the lines trough the point $y$. Gounelas and Kouvidakis studied the $g^{1}_{3}$ of these curves coming from the rulings of the corresponding quadric and their relation with the Voisin map.
Corollary 3.19.
Let $Y$ be a general cubic fourfold. Then there exists a $\Lambda\in G$ such that $\Lambda\cap Y$ is a cubic surface of type $A_{1}A_{5}$.
Proof.
Given a pair of lines $\ell_{1}$ and $\ell_{2}$ both of type I, with same residual and such that one is triple, then these are naturally contained in the cubic surface $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle\cap Y$, which is either of type $A_{1}A_{5}$ or ${\widetilde{E}}_{6}$. To discern the 2 cases, we reason as follows. A triple line through the point $y\in Y$ corresponds to a ramification point of the $g^{1}_{3}$ of the curves $C_{y}$ [GK22, Lemma 3.10]. As the general triple ramification point of one of the possibly two $g^{1}_{3}$ of the curve $C_{y}$ occurs when $C_{y}$ is smooth, we deduce that the general $\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$ cuts $Y$ in a cubic surface of type $A_{1}A_{5}$. ∎
@@
    For any $\ell\in S_{\operatorname{II}}$ the intersection $E_{\ell}\cap Y$ is a non-normal singular cubic surface.
@@
    If $\ell\in V\cap S_{\operatorname{II}}$, then $E_{\ell}\cap Y$ is of type $X_{7}$ or worse.
@@
    If $\ell\in(V\cap S_{\operatorname{II}})_{\operatorname{sing}}$ is a node, then $E_{\ell}\cap Y$ is of type $X_{8}$ or worse. In particular, there are 3780 projective spaces of dimension 3 intersecting $Y$ in a singular surface of type $X_{8}$.
@@
    For any $\ell\in S_{\operatorname{II}}$ the surface $E_{\ell}\cap Y$ is not of type $X_{9}$.
@@
Given a line $L$ of type $\operatorname{II}$, the cubic surface $E_{\ell}\cap Y$ is singular along the entire $L$. By the study of Gounelas and Kouvidakis a line as in item (2) corresponds to a triple type two line, and one as in item (3) corresponds to a type two line with two triple tangent 2-planes [GK22, comments right above Remark 4.20]. The distinction in cases comes from an elementary analysis of lines on non-normal cubic surfaces (see Appendix 5). The second statement in item (3) follows from [GK22, Theorem B].
For the last item, let $L$ be a line of type $\operatorname{II}$ such that $E_{L}\cap Y$ is a singular cubic surface of type $X_{9}$, that is a cone over a cuspidal cubic curve. Then the lines passing trough the vertex of the cone form a curve, which contains a cuspidal cubic. This is in contradiction with [GK22, Proposition 3.5]. ∎
@@ 4. The monodromy group of $\psi$ is maximal
In this section we prove that the monodromy group of ${\widehat{\psi}}$ is maximal. In order to do so, we use that the ramification of ${\widehat{\psi}}$ is simple, as proven in Section 2 and the results of Section 3 about the variety $P$.
Given a generically finite dominant morphism $f\colon X\to Y$ of degree $d$ between irreducible varieties (necessarily of the same dimension), we obtain a degree d extension of function fields $k(X)/k(Y)$, and taking the Galois closure $K/k(Y)$ of this extension, we denote by $\mathrm{Gal}_{f}=\mathrm{Gal}(K/k(Y))$. This agrees with the usual monodromy group $\mathrm{Mon}_{f}$ (see [Har79, SY21]), which is defined as the image in $S_{d}$ of the group of deck transformations of the unramified (i.e., topological) cover $X\setminus f^{-1}(\mathrm{Branch}(f))\to U$, where $U:=Y\setminus\mathrm{Branch}(f)$. Recall the following classical results for $f\colon X\to Y$ and $U$ as above.
@@
Let $X^{[s]}_{U}$ be the complement of the big diagonal in the fibre product of $X_{U}$ $s$-times with itself over $U$. In other words, the fiber of the natural morphism $X^{[s]}_{U}\to U$ consists of $s$ distinct points in the fibre of $f$.
@@
$X^{(s)}_{U}$ is irreducible if and only if $\mathrm{Mon}_{f}$ is an s-transitive subgroup of $S_{d}$.
Thanks to Corollary 2.7 we can readily apply Proposition 4.1 and deduce that the monodromy group $\mathrm{Mon}_{{\widehat{\psi}}}$ contains a transposition. In order to apply Lemma 4.2 we are led to study the irreducibility of ${\widehat{F}}^{(2)}_{U}$ which is strictly related to the variety $P$, we introduced in the previous section. The rest of the section is then devoted to prove the following which is the crucial ingredient to conclude the maximality of monodromy.
@@
The monodromy of ${\widehat{\psi}}$ is maximal, i.e. $\operatorname{Mon}_{{\widehat{\psi}}}=S_{16}$.
@@
From Corollary 2.7 and Proposition 4.1 $\operatorname{Mon}_{{\widehat{\psi}}}$ contains a transposition and thanks to Proposition 4.3 and 4.2 the action of $\operatorname{Mon}_{{\widehat{\psi}}}$ is $2$-transitive. Thus $\operatorname{Mon}_{{\widehat{\psi}}}$ contains all transpositions: Indeed up to renumbering we can assume the transposition $(1,2)\in\operatorname{Mon}_{{\widehat{\psi}}}$. By 2-transitivity there exists $\sigma\in\operatorname{Mon}_{{\widehat{\psi}}}$ such that $\sigma(1)=a$ and $\sigma(2)=b$ for any $a$ and $b$ and we have
@@
In order to prove the irreducibility of $P$ we take advantage of the map $\varphi\colon P\dashrightarrow W$ and use its various properties we have showed in Section 3, to compute the ramification of $\varphi$ at some special points. Clearly there is an open subset $U_{et}\subset W$ for which the base change $P_{U_{et}}\to U_{et}$ is étale of degree 6 and proper, but as we are interested in the ramification of $\varphi$ we need to study a slightly bigger open subset, thus we provide the following technical proposition.
@@
    The open set $U_{\mathrm{flat}}$ contains a point in $w\in W$ such that the 3-dimensional linear space $a(w)$ intersects $Y$ in a cubic surface of type $A_{1}A_{5}$.
@@
${\widehat{P}}\setminus{\widehat{\varphi}}^{-1}\left({\widehat{\varphi}}\left(E_{1}\cup E_{\pi}\right)\right)\to Y\setminus{\widehat{\varphi}}(E_{1}\cup E_{\pi})=:U$
is proper, as base change of the proper map ${\widehat{\varphi}}$, and with finite fibers, as we removed the exceptional divisor $E_{1}$). Thus it is a finite morphism. Having removed $E_{\pi}$ the source space is actually a subset of $P$. Hence, we have proven that the restriction of $\varphi$ to $P_{U}\to U$ is a finite morphism.
Moreover, the further restriction of $\varphi$ to a regular locus of $P_{U}$ is a finite morphism of non singular varieties, hence flat by [Liu02, Chapter 4, Remark 3.11]. This open set satisfies then point (1) of the claim.
${P}$${W}$${G.}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi_{G}}$$\scriptstyle{a}$
@@
${\widehat{P}}:=\mathrm{closure}\{((\ell_{1},\ell_{2}),w)\in P\times W\mid\varphi(\ell_{1},\ell_{2})=w\};$
${\widehat{P}}_{G}:=\mathrm{closure}\{((\ell_{1},\ell_{2}),g)\in P\times G\mid\varphi_{G}(\ell_{1},\ell_{2})=g\}.$
@@
(5) ${\widehat{P}}\longrightarrow{\widehat{P}}_{G},\quad\left(\ell_{1},\ell_{2},w\right)\longmapsto\left(\ell_{1},\ell_{2},a(w)\right).$
@@
By Corollary 3.19 there exists a point $w\in W$ such that $a(w)\in G$ is a point that intersects $Y$ in a surface $S$ with singularities $A_{1}A_{5}$. By Proposition 3.14 $S$ contains exactly 2 lines $\ell_{1}$ and $\ell_{2}$, that by generality are of type $I$, such that $\psi(\ell_{1})=\ell_{2}=\psi(\ell_{2})$. By the very definition of ${\widehat{P}}_{G}$, its points consist of a pair of lines and a 3-dimensional linear subspace, which contains the 2 lines, hence
${\widehat{\varphi}}_{G}^{-1}\big{(}a(w)\big{)}=\left\{\Big{(}\big{(}\ell_{1},\ell_{2}\big{)},a(w)\Big{)},\ \Big{(}\big{(}\ell_{2},\ell_{1}\big{)},a(w)\Big{)}\right\}.$
Now, as there is just one linear system of arithmetically Cohen-Macaulay generalized twisted cubics on a cubic surface of type $A_{1}A_{5}$ by Proposition 3.15 we get
${\widehat{\varphi}}^{-1}\big{(}w\big{)}=\left\{\Big{(}\big{(}\ell_{1},\ell_{2}\big{)},w\Big{)},\ \Big{(}\big{(}\ell_{2},\ell_{1}\big{)},w\Big{)}\right\}.$
@@
Notice that $\varphi_{G}|_{P}:P\dashrightarrow G$ is defined by the association $(\ell_{1},\ell_{2})\mapsto\langle L_{1},L_{2}\rangle=\langle L_{1},L_{2},R\rangle$ where $r=\psi(\ell_{1})=\psi(\ell_{2})$. Moreover, if both $\ell_{1}$ and $\ell_{2}$ are of type I, then we have $\varphi_{G}((\ell_{1},\ell_{2}))=\langle E_{\ell_{1}},E_{\ell_{2}}\rangle$, hence pairs of distinct lines of type I in $P$ are not in the indeterminacy locus of $\varphi$. Since $Z_{aCM}(S)$ is a singleton for a cubic surface $S$ of type $A_{1}A_{5}$ by Lemma 3.15, the map $\varphi$ is defined at the point $(\ell_{1},\ell_{2})$, thus we conclude that $w\not\in{\widehat{\varphi}}(E_{\pi})$.
Now we are left to prove flatness of $\varphi$ at the points $(\ell_{1},\ell_{2})$ and $(\ell_{2},\ell_{1})$ in $P$. To do so we observe that $(\ell_{1},\ell_{2})$ is a regular point of $P$ as it follows from the lemma below. ∎
@@
The singular locus of ${\widehat{F}}\times_{{\widehat{\psi}}}{\widehat{F}}$ is contained in $E\times_{{\widehat{\psi}}}E$.
@@
Let $(f_{1},f_{2})\in{\widehat{F}}\times_{{\widehat{\psi}}}{\widehat{F}}$ be a point with ${\widehat{\psi}}(f_{1})={\widehat{\psi}}(f_{2})=r$. Then its tangent space is the fiber product over the diagram
[…]
where $T_{f_{1}}{\widehat{F}}$ and $T_{f_{2}}{\widehat{F}}$ have dimension 4 as ${\widehat{F}}$ is a non-singular fourfold. If one among $f_{1}$ and $f_{2}$ is not in the ramification of ${\widehat{\psi}}$, which coincide with the exceptional divisor $E$, then $d{\widehat{\psi}}$ is an isomorphism at that point and the fiber product is a vector space of dimension 4, hence the claim. ∎
@@
$(\ell_{1},\ell_{2})\quad\text{and}\quad(\ell_{2},\ell_{1}).$
Using the involution $\sigma$ on $F\times F$ switching the factors, which clearly restrict to an involution of $P$, we see that $P$ at $(\ell_{1},\ell_{2})$ is locally isomorphic to $P$ at $\sigma((\ell_{1},\ell_{2}))=(\ell_{2},\ell_{1})$. As $\varphi$ is flat, proper and of degree 6 at the point $q$, we conclude that the schematic preimage $\varphi^{-1}(q)$ is a finite scheme with multiplicity 3 both at $(\ell_{1},\ell_{2})$ and $(\ell_{2},\ell_{1})$. We pick now a point $p\in U_{\mathrm{et}}$ close enough to $q$ and focus on the preimages $\varphi^{-1}(p)=\{p_{1},\cdots,p_{6}\}$: If we let $p$ tend to $q$, then 3 points out of the 6 in $\varphi^{-1}(p)$, say $p_{1},p_{2}$ and $p_{3}$, tend to $(\ell_{1},\ell_{2})$ and the other 3 points, i.e. $p_{4},p_{5}$ and $p_{6}$, tend to $(\ell_{2},\ell_{1})$. Moreover, $\sigma$ maps the first triple to the second one, i.e. $\sigma(\{p_{1},p_{2},p_{3}\})=\{p_{4},p_{5},p_{6}\}$, thus $p_{1},p_{2}$ and $p_{3}$ lie in the same connected component of $P$, in other words the quotient of $P$ by the switching involution is path-connected. By Lemma 4.6 we deduce that it is irreducible too.
Hence, it is now sufficient to show that for some point $p_{1}\in P_{U_{et}}$, the points $p_{1}$ and $\sigma(p_{1})$ lie in the same connected component of $P_{U_{et}}$. To do so, we consider one transposition $\tau$ in $\operatorname{Mon}_{\psi}$, whose existence is granted by Proposition 4.1 and Corollary 2.7. If we consider it as an element of deck transformations, this means that there exists a point $r\in F\setminus\operatorname{Branch}({\widehat{\psi}})$, and a loop $\gamma$ centered at $r$ such that $\gamma$ acts on the fiber of $r$ exchanging two points, say $\ell_{1}\mapsto\ell_{2}$ and $\ell_{2}\mapsto\ell_{1}$, and as identity on the remaining other points of the fiber. We now consider the action under deck transformation of $\gamma$ for the natural map $P\dashrightarrow F$ defined by $\psi$. A moment’s thought tells us that under the action of $\gamma$ the point $p=(\ell_{1},\ell_{2})\in P$, with $\psi(\ell_{1})=\psi(\ell_{2})=r$, is mapped to $(\ell_{2},\ell_{1})$. Thus $P$ is irreducible as we claimed. ∎
5. Appendix
In this appendix we study singular cubic surfaces that are degenerations of a cubic surface with exactly 4 singular points of type $A_{1}$. These surfaces are called Cayley.
The stratification of the space of cubic surfaces given by the type of singularities has been studied by Bruce in [Bru80]. From his analysis Cayley cubics can degenerate to a singular surface of type $2A_{1}A_{3}$, which in turn degenerates to a singular surface $A_{1}A_{5}$, which in turn degenerates to surfaces with simple elliptic singularities and then to non-normal ones. Here is the diagram of degenerations:
$4A_{1}\succeq 2A_{1}A_{3}\succeq A_{1}A_{5}\succeq{\widetilde{E}}_{6}\succeq X_{6}\succeq X_{7}\succeq X_{8}\succeq X_{9}$
where ${\widetilde{E}}_{6}$ denotes the cubics with one simple elliptic singularity and $X_{i}$ are non-normal cubic surfaces. We discuss them in details in the subsequent subsections, focussing on lines and generalised twisted cubics and the restrictions of the Voisin maps $\varphi$, and $\psi$ to a cubic surface.

### v2

@@ 3.13.1. Simple elliptic cubic surfaces
  • the residual to the line $p\in E\simeq F(S)$ is given by 2p-2p;
@@
We explore now some connections of our study of singular cubic surfaces with results of Gounelas and Kouvidakis [GK22]. As a consequence, we deduce the existence of a surface of type $A_{1}A_{5}$, which is a fundamental little ingredient for the main theorem.
Let $Y$ be a smooth cubic fourfold. The natural map $\gamma\colon\mathbb{L}\to Y$ from the universal family of lines to the cubic fourfold is a fibration in (2,3)-complete intersection curves. Given a point $y\in Y$, its fibre $\gamma^{-1}(y)$ consists of the lines trough the point $y$. Gounelas and Kouvidakis studied the $g^{1}_{3}$ of these curves coming from the rulings of the corresponding quadric and their relation with the Voisin map. They proved that for the general cubic fourfolds the surface $V$ of triple lines intersects $S_{\operatorname{II}}$ in an irreducible curve with 3780 nodes. In particular, the general triple line is of type II [GK22, Theorem B].
@@
    For any $L\in S_{\operatorname{II}}$ the intersection $\Lambda_{L}\cap Y$ is a non-normal singular cubic surface.
@@
    If $L\in V\cap S_{\operatorname{II}}$, then $\Lambda_{L}\cap Y$ is of type $X_{7}$ or worse.
@@
    If $L\in(V\cap S_{\operatorname{II}})_{\operatorname{sing}}$ is a node, then $\Lambda_{L}\cap Y$ is of type $X_{8}$ or worse. In particular, there are 3780 projective spaces of dimension 3 intersecting $Y$ in a singular surface of type $X_{8}$.
@@
    For any $L\in S_{\operatorname{II}}$ the surface $\Lambda_{L}\cap Y$ is not of type $X_{9}$.
@@
Given a line $L$ of type $\operatorname{II}$, the cubic surface $\Lambda_{L}\cap Y$ is singular along the entire $L$. By the study of Gounelas and Kouvidakis a line as in item (2) corresponds to a triple line of type II, and one as in item (3) corresponds to a type two line with two triple tangent 2-planes [GK22, comments right above Remark 4.20]. The distinction in cases comes from an elementary analysis of lines on non-normal cubic surfaces (see Appendix A). The second statement in item (3) follows from [GK22, Theorem B].
For the last item, let $L$ be a line of type $\operatorname{II}$ such that $\Lambda_{L}\cap Y$ is a singular cubic surface of type $X_{9}$, that is a cone over a cuspidal cubic curve. Then the lines passing trough the vertex of the cone form a curve, which contains a cuspidal cubic. This is in contradiction with [GK22, Proposition 3.5]. ∎
Proposition 3.16.
Let $Y$ be a general cubic fourfold, and let $L_{1}$ and $L_{2}$ be general distinct lines such that $L_{2}$ is residual to $L_{1}$, and $L_{2}$ is of type II and triple. Then $\langle{L_{1}},\Lambda_{L_{2}}\rangle$ is a 3-dimensional space which cuts $Y$ in a cubic surface of type either $A_{1}A_{5}$, or ${\widetilde{E}}_{6}$.
Proof.
By Lemma 3.4 a general point of $P$ consists of a pair of lines whose linear span intersects $Y$ in a Cayley cubic. Thus, if $L_{1}$ and $L_{2}$ are as in the hypothesis, then the 3 dimensional linear space $\langle L_{1},\Lambda_{L_{2}}\rangle$ intersects $Y$ in a cubic surface $S$ which is a degeneration of a Cayley cubic and contains $L_{2}$ as a triple line on the surface. By the analysis of lines on cubic surfaces, $S$ is either of type $A_{1}A_{5}$, ${\widetilde{E}}_{6}$ or non-normal. As both $V$ and $S_{\operatorname{II}}$ are surfaces, both the families of non-normal cubic surfaces and of surfaces containing a triple line, which is triple on the surface, have dimension 2. By Corollary 3.15 and the Appendix the general non-normal cubic surface (which is a linear section of $Y$) does not contain a triple line, thus the general surface containing a triple line is of type $\widetilde{E}_{6}$ or $A_{1}A_{5}$. ∎
Corollary 3.17.
Let $Y$ be a general cubic fourfold. Then there exists $\Lambda\in G$ such that $\Lambda\cap Y$ is a cubic surface of type $A_{1}A_{5}$.
Proof.
Let $L_{1},L_{2}$ be general distinct lines such that $L_{2}$ is a triple line of type II and $L_{2}$ is residual to $L_{1}$. We consider the projective space $\Lambda:=\langle\Lambda_{L_{2}},L_{1}\rangle$ and the singular cubic surface $S:=\Lambda\cap Y$ it defines. In particular, the line $L_{2}$ is triple on $S$ and is residual to $L_{1}$. By Proposition 3.16 the surface $S$ is either singular of type $A_{1}A_{5}$ or isomorphic to a cone over an elliptic curve.
In other words, the association $(L,R)\mapsto\Lambda$ for pairs of lines as above defines a rational map $\widehat{\psi}^{-1}(V)\dashrightarrow G$ whose image lies in the union of loci $G_{A_{1}A_{5}}\cup G_{\widetilde{E}_{6}}$. This is generically finite by Propositions 3.14 and 3.12. Since $G_{\widetilde{E}_{6}}$ is 1-dimensional by the discussion in [NO21, Section 1.4] the locus $G_{A_{1}A_{5}}$ is not empty. ∎
@@ 4. The monodromy group of $\psi$ is maximal
In this section we prove that the monodromy group of $^{\widehat{\psi}}$ is maximal. In order to do so, we use that the ramification of $^{\widehat{\psi}}$ is simple, as proven in Section 2 and the results of Section 3 about the variety $P$.
Given a generically finite dominant morphism $f\colon X\to Y$ of degree $d$ between irreducible varieties (necessarily of the same dimension), we obtain a degree d extension of function fields $k(X)/k(Y)$, and taking the Galois closure $K/k(Y)$ of this extension, we denote by $\mathrm{Gal}_{f}=\mathrm{Gal}(K/k(Y))$. This agrees with the usual monodromy group $\mathrm{Mon}_{f}$ (see [Har79, SY21]), which is defined as the image in the symmetric group $S_{d}$ of the group of deck transformations of the unramified (i.e., topological) cover $X\setminus f^{-1}(\mathrm{Branch}(f))\to U$, where $U:=Y\setminus\mathrm{Branch}(f)$. Recall the following classical results for $f\colon X\to Y$ and $U$ as above.
@@
Let $X^{(s)}_{U}$ be the complement of the big diagonal in the fibre product of $X_{U}$ $s$-times with itself over $U$. In other words, the fibre of the natural morphism $X^{(s)}_{U}\to U$ consists of $s$ distinct points in the fibre of $f$.
@@
The variety $X^{(s)}_{U}$ is irreducible if and only if $\mathrm{Mon}_{f}$ is an s-transitive subgroup of $S_{d}$.
We want to apply now these results to the resolution of the Voisin map $\widehat{\psi}\colon\widehat{F}\to F$. For the remainder of the section, let $Y$ be a general cubic fourfold, $F$ be its Fano variety of lines and $U\subset F$ be the complement of the branch locus of $\widehat{\psi}$. We recall that $P\subset F\times F$ introduced in §3.3 is mapped onto $W$ under the map $\varphi\colon F\times F\dashrightarrow Z$. Thanks to Proposition 2.7 we can readily apply Proposition 4.1 and deduce that the monodromy group $\mathrm{Mon}_{{\widehat{\psi}}}$ contains a transposition. In order to apply Lemma 4.2 we are led to study the irreducibility of ${\widehat{F}}^{(2)}_{U}$ which is strictly related to the variety $P$. The rest of the section is then devoted to prove the following which is the crucial ingredient to conclude the maximality of monodromy.
@@
The monodromy of $^{\widehat{\psi}}$ is maximal, i.e. $\operatorname{Mon}_{{\widehat{\psi}}}=S_{16}$.
@@
From Propositions 2.7 and 4.1 $\operatorname{Mon}_{{\widehat{\psi}}}$ contains a transposition and thanks to Proposition 4.3 and 4.2 the action of $\operatorname{Mon}_{{\widehat{\psi}}}$ is $2$-transitive. Thus $\operatorname{Mon}_{{\widehat{\psi}}}$ contains all transpositions: Indeed up to renumbering we can assume the transposition $(1,2)\in\operatorname{Mon}_{{\widehat{\psi}}}$. By 2-transitivity there exists $\sigma\in\operatorname{Mon}_{{\widehat{\psi}}}$ such that $\sigma(1)=a$ and $\sigma(2)=b$ for any $a$ and $b$ and we have
@@
In order to prove the irreducibility of $P$ we take advantage of the map $\varphi\colon P\dashrightarrow W$ and use its various properties we have showed in Section 3, to compute the ramification of $\varphi$ at some special points. Clearly there is an open subset $U_{et}\subset W$ for which the base change $P_{U_{et}}\to U_{et}$ is étale of degree 6 and proper, but as we are interested in the ramification of $\varphi$ we need to study a slightly bigger open subset, thus we provide the following technical proposition. Recall that $Z$ admits a generically finite map $a\colon Z\dashrightarrow G$ whose indeterminacy locus is $Y$.
@@
    There exists a point $w\in U_{\mathrm{flat}}$ such that the 3-dimensional linear space $a(w)$ intersects $Y$ in a cubic surface of type $A_{1}A_{5}$.
@@
Let $P\xleftarrow{\pi}{\widehat{P}}\xrightarrow{{\widehat{\varphi}}}W$ be any resolution of indeterminacy of $\varphi$ and ${\widehat{P}}\xrightarrow{\varphi_{1}}P_{\mathrm{Stein}}\xrightarrow{\varphi_{2}}W$ be its Stein factorization. We denote by $E_{\pi}$ and $E_{1}$ the exceptional loci of $\pi$ and $\varphi_{1}$ respectively. Then the morphism
${\widehat{P}}\setminus{\widehat{\varphi}}^{-1}\left({\widehat{\varphi}}\left(E_{1}\cup E_{\pi}\right)\right)\to W\setminus{\widehat{\varphi}}(E_{1}\cup E_{\pi})=:U$
is proper, as it is the base change of the proper morphism $^{\widehat{\varphi}}$. It has finite fibres, as we removed the exceptional divisor $E_{1}$. Thus it is a finite morphism. Having removed $E_{\pi}$, the source space is actually a subset of $P$. Hence, we have proven that the restriction of $\varphi$ to $P_{U}\to U$ is a finite morphism.
Moreover, the further restriction of $\varphi$ to $U\setminus\varphi(P_{\mathrm{sing}})$ is a finite morphism of non singular varieties, hence flat by [Liu02, Chapter 4, Remark 3.11]. This open set satisfies point (1) of the claim.
@@
${\widehat{P}}:=\mathrm{closure}\{((L_{1},L_{2}),w)\in P\times W\mid\varphi(L_{1},L_{2})=w\};$
${\widehat{P}}_{G}:=\mathrm{closure}\{((L_{1},L_{2}),g)\in P\times G\mid\varphi_{G}(L_{1},L_{2})=g\}.$
@@
(12) ${\widehat{P}}\longrightarrow{\widehat{P}}_{G},\quad\left(L_{1},L_{2},w\right)\longmapsto\left(L_{1},L_{2},a(w)\right).$
@@
By Corollary 3.17 there exists a point $w\in W$ such that $a(w)\in G$ is a point that intersects $Y$ in a surface $S$ with singularities $A_{1}A_{5}$. By Proposition 3.12 the surface $S$ contains exactly 2 lines $L_{1}$ and $L_{2}$, that by generality are of type II, such that $\psi(L_{1})=L_{2}=\psi(L_{2})$. By the very definition of ${\widehat{P}}_{G}$, its points consist of a pair of lines and a 3-dimensional linear subspace, which contains the 2 lines, hence
${\widehat{\varphi}}_{G}^{-1}\big(a(w)\big)=\left\{\Big(\big(L_{1},L_{2}\big),a(w)\Big),\ \Big(\big(L_{2},L_{1}\big),a(w)\Big)\right\}.$
Now, as there is just one linear system of arithmetically Cohen-Macaulay generalised twisted cubics on a cubic surface of type $A_{1}A_{5}$ by Lemma 3.13 we get
${\widehat{\varphi}}^{-1}\big(w\big)=\left\{\Big(\big(L_{1},L_{2}\big),w\Big),\ \Big(\big(L_{2},L_{1}\big),w\Big)\right\}.$
@@
Notice that $\varphi_{G}|_{P}:P\dashrightarrow G$ is defined by the association $(L_{1},L_{2})\mapsto\langle L_{1},L_{2}\rangle=\langle L_{1},L_{2},R\rangle$ where $R=\psi(L_{1})=\psi(L_{2})$. Moreover, if both $L_{1}$ and $L_{2}$ are of type I, then we have $\varphi_{G}((L_{1},L_{2}))=\langle\Lambda_{L_{1}},\Lambda_{L_{2}}\rangle$, hence pairs of distinct lines of type I in $P$ are not in the indeterminacy locus of $\varphi$. Since $Z_{aCM}(S)$ is a singleton for a cubic surface $S$ of type $A_{1}A_{5}$ by Lemma 3.13, the map $\varphi$ is defined at the point $(L_{1},L_{2})$, thus we conclude that $w\not\in{\widehat{\varphi}}(E_{\pi})$.
Now we are left to prove that $(L_{1},L_{2})$ is a regular point of $P$, so that $\varphi$ is flat at the points $(L_{1},L_{2})$ and $(L_{2},L_{1})$ in $P$. We prove the regularity of $P$ at the point $(L_{1},L_{2})$ in the lemma below. ∎
@@
The singular locus of $^{\widehat{F}}\times_{{\widehat{\psi}}}{\widehat{F}}$ is contained in $E\times_{{\widehat{\psi}}}E$.
@@
Let $(f_{1},f_{2})\in{\widehat{F}}\times_{{\widehat{\psi}}}{\widehat{F}}$ be a point with ${\widehat{\psi}}(f_{1})={\widehat{\psi}}(f_{2})=R$. Then its tangent space is the fibre product over the diagram
$^{T_{f_{1}}{\widehat{F}}}$$^{T_{f_{2}}{\widehat{F}}}$$T_{R}F$$\scriptstyle{d{\widehat{\psi}}}$$\scriptstyle{d{\widehat{\psi}}}$
where $T_{f_{1}}{\widehat{F}}$ and $T_{f_{2}}{\widehat{F}}$ have dimension 4 as $^{\widehat{F}}$ is a non-singular fourfold. If one among $f_{1}$ and $f_{2}$ is not in the ramification of $^{\widehat{\psi}}$, which coincides with the exceptional divisor $E$, then $d{\widehat{\psi}}$ is an isomorphism at that point and the fibre product is a vector space of dimension 4, hence the claim. ∎
@@
$(L_{1},L_{2})\quad\text{and}\quad(L_{2},L_{1}).$
Using the involution $\sigma$ on $F\times F$ switching the factors, which clearly restrict to an involution of $P$, we see that $P$ at $(L_{1},L_{2})$ is locally isomorphic to $P$ at $\sigma((L_{1},L_{2}))=(L_{2},L_{1})$. As $\varphi$ is flat, proper and of degree 6 at the point $q$, we conclude that the schematic preimage $\varphi^{-1}(q)$ is a finite scheme with multiplicity 3 both at $(L_{1},L_{2})$ and $(L_{2},L_{1})$. We pick now a point $p\in U_{\mathrm{et}}$ close enough to $q$ and focus on the preimages $\varphi^{-1}(p)=\{p_{1},\cdots,p_{6}\}$: If we let $p$ tend to $q$, then 3 points out of the 6 in $\varphi^{-1}(p)$, say $p_{1},p_{2}$ and $p_{3}$, tend to $(L_{1},L_{2})$ and the other 3 points, i.e. $p_{4},p_{5}$ and $p_{6}$, tend to $(L_{2},L_{1})$. Moreover, $\sigma$ maps the first triple to the second one, i.e. $\sigma(\{p_{1},p_{2},p_{3}\})=\{p_{4},p_{5},p_{6}\}$, thus $p_{1},p_{2}$ and $p_{3}$ lie in the same connected component of $P$, in other words the quotient of $P$ by the switching involution is path-connected. By Lemma 4.6 we deduce that $P/\sigma$ is irreducible.
[…]
Appendix A Singular cubic surfaces
In this appendix we treat degenerations of Cayley cubics we introduced in Diagram (10), focussing on lines, generalised twisted cubics and the restrictions of the Voisin maps $\varphi$, and $\psi$ to a cubic surface.